OpenStudy (anonymous):

Maximize P = xy with x + y = 6.

OpenStudy (ryan123345):

Hey there! I see that you are new to openstudy. This is a great website on getting help with questions of any subject. It is a fun way of learning! I wish you good luck on openstudy and dont forget to read the http://openstudy.com/code-of-conduct so you know the rules and regulations :) Hope you have fun!

OpenStudy (anonymous):

P = x(6-x) = -x^2 + 6x Now use first derivative test to differentiate P w.r.t x and find the critical points

OpenStudy (anonymous):

Without calculus, you can find the vertex of the parabola by using x=-b/2a with a=coefficient of square term and b=coefficient of x to first power term. Once you find x, substitute it back into your equation to find y, and this will be your maximum value.